OpenStudy (anonymous):

What is the name for the rocky upper layer of the solid earth? A. biosphere B. magnetosphere C. atmosphere D. lithosphere

OpenStudy (shadowlegendx):

What do you think the answer is?

OpenStudy (shadowlegendx):

@bayanAssad123

OpenStudy (anonymous):

a

OpenStudy (anonymous):

or c

OpenStudy (uh60blackhawk):

no that is incorrect atmosphere is not a rock.

OpenStudy (anonymous):

oh well thanks can u help me with a math problem

OpenStudy (uh60blackhawk):

it should be the lithoshere

OpenStudy (anonymous):

yea i got it i need help with math

OpenStudy (uh60blackhawk):

oh ok tag me might be able to hellp

OpenStudy (anonymous):

im knew how do u tag lol

OpenStudy (anonymous):

Which variable expression matches the word phrase? Column A Column B 1. the sum of 10 and the product of 7 and a number 2. 10 more than the product of 7 and a number 3. the difference of 10 and the product of 7 and a number 4. 10 more than the quotient of a number and 7 A. 10 – 7y B. 10 + (y ÷ 7) C. 10 + 7y

OpenStudy (uh60blackhawk):

Sorry can't remember how to do these.

OpenStudy (anonymous):

oh well ty anyways :/

OpenStudy (shadowlegendx):

This is simple. Translating words into expressions and/or equations

OpenStudy (shadowlegendx):

Let me give you the keywords. Sum is the total amount from the adding of two or more numbers For example: 1 + 2 = 3 3 is the sum a number basically means that this number is a variable For example: 5 - x = ? x is the "number" Product is the result of multiplication For example: 7 x 7 = 49 49 is the product More than basically means addition of two numbers More than Two more than two 2 + 2 + = more than Quotient is the result of division For Example: \[6 \div 3 = 2 \] 2 is the quotient

OpenStudy (shadowlegendx):

@bayanAssad123

OpenStudy (shadowlegendx):

the difference is the subtraction of one number from another For example: The difference of 3 and 1 3 - 1

OpenStudy (shadowlegendx):

@bayanAssad123 Use this to help solve your problem. If you need further help, let me know :)